likeabossssssss:

In the waiting room of a vet's office, there are 4 cats, 5 dogs, 1 lizard, and 1 hamster. A represents the event that a randomly selected animal in this waiting room is a dog. What is the probability of the complement of event A? Enter your answer as a fraction, in simplified form, in the box.

likeabossssssss:

@KjSaif @SkyVoltage43

SkyVoltage43:

imma pull a @Justin234 and copy from brainly The answer would be p(dog)=5/11 because you need to find the total amount of animals in the waiting room. You then would take however many dogs there are and make that into a probability

Justin234:

Thats amazing news im glad you could be me to the answer <3

KjSaif:

The complement of event A is not A and p(A)+p(notA)=1 since the categories are mutual exclusive, i.e. a dog is not a cat. So p(not A)=1-p(A)=1-5/11=6/11.

SkyVoltage43:

go with his answer^

Justin234:

OHHHHH I FOUND THE ANSWER The answer would be p(dog)=5/11 because you need to find the total amount of animals in the waiting room. You then would take however many dogs there are and make that into a probability

likeabossssssss:

ffs

Justin234:

XD oops now im copiing from Sky and Brainly... My bad

likeabossssssss:

so its 6/11

Justin234:

yes

likeabossssssss:

k